# Modeling and Optimizing User Preferences in AI Copilots: A Comprehensive Survey and Taxonomy

### v1

@@ Abstract
AI copilots, context-aware, AI-powered systems designed to assist users in tasks such as software development and content creation, are becoming integral to modern workflows. As these systems grow in capability and adoption, personalization has emerged as a cornerstone for ensuring usability, trust, and productivity. Central to this personalization is preference optimization: the ability of AI copilots to detect, interpret, and align with individual user preferences. While personalization techniques are well-established in domains like recommender systems and dialogue agents, their adaptation to interactive, real-time systems like AI copilots remains fragmented and underexplored. This survey addresses this gap by synthesizing research on how user preferences are captured, modeled, and refined within the design of AI copilots. We introduce a unified definition of AI copilots and propose a phase-based taxonomy of preference optimization strategies, structured around pre-interaction, mid-interaction, and post-interaction stages. We analyze techniques for acquiring preference signals, modeling user intent, and integrating feedback loops, highlighting both established approaches and recent innovations. By bridging insights from AI personalization, human-AI collaboration, and large language model adaptation, this survey provides a structured foundation for designing adaptive, preference-aware AI copilots. It offers a holistic view of the available preference resources, how they can be leveraged, and which technical approaches are most suited to each stage of system design.
@@ 1 Introduction
AI-powered, context-aware systems—commonly referred to as AI copilots—are rapidly emerging as collaborative assistants for developers, analysts, and knowledge workers engaged in complex, high-context tasks. These systems are integrated into user workflows and designed to understand evolving goals, adapt to individual preferences, and provide real-time, personalized support. However, achieving meaningful personalization in such systems presents unique challenges. User intent is often ambiguous or changes over time, tasks may span multiple domains, and responses must be dynamically adapted to nuanced contexts. These systems must be capable of aligning with dynamic user preferences, as personalization is a key success factor for digital assistants, enabling them to adapt their responses and behavior to individual users’ routines and context [1].
At the core of this personalization lies preference optimization—the ability of a system to detect, interpret, and align with user preferences [2]. While advances in preference modeling have gained traction in machine learning and recommender systems, their integration into real-time, interactive systems like AI copilots remains limited and fragmented. Designing copilots that can continuously adapt to evolving preferences in high-context environments is still a major open challenge.
@@
Research related to this survey broadly falls into two categories: (i) AI systems and intelligent assistants, and (ii) preference optimization techniques. While each has been explored independently in the literature, their integration remains underexamined. The evolution of virtual assistants such as Google Assistant, Siri, Cortana, and Alexa has been well documented, with studies focusing on their core functionalities, user experience, and limitations in areas such as voice recognition, contextual awareness, and multi-turn dialogue management [3]. In enterprise contexts, digital assistants have been examined through the lens of business integration, workflow support, and human-computer collaboration, highlighting both their organizational potential and associated risks [1]. Concerns around privacy and security in these systems have also been widely addressed, particularly in relation to data handling and unauthorized access [4]. Although the design and evaluation of AI copilots in retail has been recently analyzed through Microsoft’s case studies, the work is limited in scope and does not offer a comprehensive perspective on the broader cognitive assistant landscape [5]. On the other hand, preference optimization—especially in the context of aligning AI behavior with user intent—has seen significant theoretical and practical development. A recent comprehensive survey on Direct Preference Optimization (DPO) presents an in-depth analysis of its theoretical foundations, training methodologies, benchmark datasets, and diverse application scenarios [2]. However, such works primarily examine preference modeling in isolation and do not address its integration within interactive systems like AI copilots or digital assistants.
While existing studies contribute significantly to our understanding of intelligent assistant technologies and preference optimization methods, their **integration in the context of AI copilots remains limited and underexplored**. As a result, there is no consolidated view of how preference signals can be effectively incorporated into the behavior of real-time, context-aware systems. This survey aims to bridge these two streams of research by examining preference optimization techniques within the design of AI copilots—systems that operate collaboratively with users in high-context, evolving environments. Specifically, we address how preferences can be detected, modeled, and refined to enhance personalization and system responsiveness. The main contributions of this work are summarized as follows:
  • A unified and literature-grounded definition of AI copilots is proposed, synthesizing diverse descriptions and terminology from recent studies across domains.
@@
We organized the content of our survey as follows: Section 2 provides foundational background by examining the historical development and evolving terminology associated with AI copilots. Section 3 introduces a unified, literature-informed definition of AI copilots, integrating insights from diverse fields. Section 4 presents a conceptual AI copilot architecture focused on preference optimization, offering a structured analysis of core components, including preference sources, detection techniques, personalization strategies, and feedback-driven refinement. Finally, Section 5 concludes with a summary of key insights and implications for future research.
@@ 2.1 AI Copilot Related Terms
2.1.1 Cognitive Assistants
One of the concepts related to the scope of our survey in the literature is AI Assistant, which shouldn’t be used interchangeably with AI Copilot. As a general definition of AI Assistant, Danaher [6] defines it as a program that tries to optimize a required target output based on the possible range of options. This notion highlights the nature of AI assistants as an optimization tool rather than as collaborative enhancers of human abilities. As a more recent idea, AI co-pilots aim to work alongside people to improve their decision-making and capacity in more complex scenarios. The AI Co-Pilot concept, as highlighted in Sellen et al. (2023)[7], is defined as a collaborative partner aiding humans, rather than just a tool. It emphasizes assisting humans with insightful support in professional and industrial settings, enhancing the human-AI partnership with its context-aware capabilities. Therefore, while AI Assistants serve the broader consumer base, AI Co-Pilots are specifically designed with developers in mind as their target user group.
Starting with Cognitive Computing, this broad term encompasses hardware or software systems designed to mimic human cognitive abilities, as explored by Kelly (2015)[8]. In such a context, Metacognition described by Schmorrow (2005)[9] as a nuance expansion, extends the paradigm by fostering a symbolic, closed-loop integration of humans and machines. In such a continuous feedback loop, the communication channel for information sharing and decision-making is done by cognitive functions. Delving into a practical application, the Joint Cognitive System, highlighted by Hollnagel (2005)[10], presents an interactive environment where goal-oriented human and machine elements collaboratively share and integrate their cognitive abilities within specific contexts. Rooted in a human-centric approach, this strategy accentuates the synergy of human creativity with AI’s analytical capabilities, ensuring adaptability across scenarios. Cognitive assistants stand at the forefront of Human-AI Collaboration, merging human insight with AI’s vast processing capacity to improve problem-solving and decision-making [11]. Through studies in hybrid intelligence systems and empathic AI-supported conversations, these technologies demonstrate their capacity to enhance teamwork, innovation, and human empathy [12, 13]. This synergy underscores the significant benefits of blending AI with human insights, positioning cognitive assistants as key facilitators of effective human-AI partnerships.
@@ 2.1.2 Digital Assistant
The term Digital Assistant (DA) is frequently interchanged with Virtual Assistant (VA), yet it represents a broader concept. In a study about its anatomy, Wells et al. [14] referred to it as a versatile tool and amorphous compound technology. Mobility, rapid data analysis capability and supporting multiple interfaces are among its key features making it ideal for task delegation and advice provision applications. The study by Knote et al. [15] highlights a key feature of Digital Assistants (DAs): their ability to merge the simplicity of intuitive natural language dialogue with personalized, context-specific assistance. This study focuses on categorizing the different forms of the Smart Personal Assistant (SPA), which is essentially a personalized Digital Assistant. Thus, the classification of SPAs into text-based systems, often termed Chatbots, and voice-based systems, as conducted by Cowan et al. [16], can be applied to DAs. By considering Chatbots as a form of Digital Assistant, all the categories of employed technologies in their construction can be categorized under DA. This allows DA’s boundaries to be expanded to cover a wider range of technologies, from simple rule-based Chatbots with limited capabilities to intelligent voice-based VAs with speech recognition capability [17].
Table 1: Comparison of AI Copilot-Related Terms and Technologies
@@
Cognitive Assistant Advanced decision support through reasoning and learning Proactive / Semi-autonomous Multimodal (text, voice, context-aware UI) Often domain-specific (e.g., healthcare, law)
@@
Chatbot Text-based conversational responses Reactive Text Often domain-specific (e.g., customer support, retail)
@@
AI Copilot Context-aware collaboration on complex or creative tasks Proactive Multimodal (text, code, UI, APIs) Task- or profession-specific (e.g., software development, writing)
@@ 2.1.3 Virtual Assistant
As a good transitional definition from DA to VA, Apple Siri and Amazon Alexa as two popular intelligent assistants referred to as voice-based DAs [1, 18]. Their ability to perform hands-free operations, context understanding and handle complex conversations are among the key features mentioned by Wellsandt et.al study [14]. Besides their speech recognition ability to understand voice commands and communicate over the same channel [4], voice-activation ability made them suitable for integrating into daily use devices [19]. Although they are mainly referred to as VAs, Smart Personal Assistant (SPA) is another term used for Apple Siri and Amazon Alexa as two device-integrated assistants focusing on their personalized assisting[15]. This positions them within the broader spectrum of Digital Assistants, primarily known for their voice-based interfaces, but also encompassing text-based interaction as an additional feature, distinguishing another subset of such assistants [20].
@@ 2.1.4 Chatbot
Chatbots are specialized programs tailored for customer service, helping users find desired information [21]. This focused definition underscores their application and emphasizes their role, setting them apart from other digital assistants despite shared technologies and capabilities. A defining feature of chatbots, as highlighted in literature [22], is simulating near-human conversations using natural language or text, to create an impression of engaging in dialogue with a human [23]. Emerging from the field of Human-Computer Interaction (HCI), which centres on enabling users to communicate with computers, Dialogue Systems is a term often used to describe Chatbots [21, 24]. Conversational Systems, Conversational Agents, and Chatter-bot are among other terms used interchangeably to refer to the same term [22]. While there are different criteria for Chatbot classification in the literature, they are primarily classified by their applications, falling into either Task-oriented or Non-task-oriented (open-ended or simple chit-chat) categories [21, 23].
Task-oriented chatbots generally consist of domain-specific hand-crafted rules with limited and focused conversation context analysis and are typically utilized in devices for convenience. Non-task-oriented or informative Chatbots, often handle random dialogue contexts without aiming for a specific goal and may require online search capabilities to operate effectively [24, 23]. Based on the Knowledge Domain, chatbots can be categorized into closed-domain or open-domain, where the open-domain modeling, a challenging field, is predominantly addressed using Neural Network (NN) approaches, including both Retrieval-based and Generative (Generation-based) Chatbots [22]. Retrieval-based chatbots support immediate response generation in e-business by ranking candidate responses based on similarity measurements between the query and knowledge base of QA pairs. Data-driven Generative Chatbots utilize Natural Language Processing (NLP) to extract content from queries and learn from entire dialogue contexts, rather than solely focusing on the current turn. These chatbots apply both supervised methods, which rely on annotated data and manual feature extraction, and unsupervised methods, which automatically learn features from unlabeled data. Responses are generated word by word using neural network techniques such as RNN, LSTM, and Deep Seq2Seq, all of which operate based on a learned vocabulary [22]. In this regard, RAG-based approaches as a more recent approach, combine the previous techniques for creating a better open-domain chatbot and even injecting more domain-specific knowledge by additional training [25].
@@ 2.1.5 Expert System
As one of the old terms in cognitive computing, an Expert System (ES) is defined as a computer program that emulates the behavior of a human expert [26]. They are a branch of AI (that is a high-level term) that is used for transferring task-specific knowledge from human to computer in the way that it continuously changes and obtains new knowledge[27]. Used in problem-oriented domains including social science, cognitive science and human behaviour analysis, ES has two main parts that enable it to produce expert-like advice. As the first main component, the knowledge base, constructed by knowledge engineers using the suitable notation of language, stores the codified collected knowledge. The inference engine, guided by a suitable reasoning method chosen to ensure efficient use of the knowledge, then uses the knowledge to develop a conclusion [26]. For transparency reasons, the system can benefit from an Explanation Facility to elucidate the logic behind the conclusion [28]. Although Rule-Based Systems (RBS) or Knowledge-Based Systems (KBS) are used to refer to the same term [29, 28], others consider them as main streams of ES[27]. In the RBS, knowledge representation is a collection of IF-THEN rules. KBS as a human-centred design, includes knowledge engineering tools and a user interface for this reason. Case-Based Reasoning (CBR) systems in one category of ESs adapt solutions to new challenges by comparing them to past cases based on set parameters. Meanwhile, Fuzzy Expert Systems (Fuzzy ESs) in another category, embrace uncertainty in human reasoning, transcending binary logic for nuanced decision-making in complex situations where traditional approaches fall short [27].
The Belief Rule-Based Expert System (BRBES) merges the accuracy of expert systems with the adaptability of data-driven models, adeptly handling quantitative and qualitative data to address complexities in uncertain environments [30]. It manages uncertainties—like fuzziness, randomness, and ignorance—through belief degrees and employs the Evidence Reasoning (ER) algorithm for seamless data integration. This approach makes BRBES a transparent and dependable solution for decision-making in situations marked by scarce or ambiguous information.
@@ 3 AI Copilot Definition
AI copilots have been described as intelligent assistants designed to support users in executing cognitive tasks [5]. Their architectures are typically composed of large language models, mechanisms for knowledge retrieval and action execution, orchestration modules to manage subtasks, structured system prompts, and responsible AI safeguards to ensure alignment with human intent. In domain-specific applications such as pathology, copilots have been tailored through fine-tuning on real-world instructions to address complex tasks like question answering [31]. These systems are often multimodal and interactive, capable of processing both visual and textual inputs, and are used to assist human-in-the-loop decision-making processes. In programming contexts, copilots have been implemented to provide real-time suggestions that support knowledge workers and improve productivity, while allowing users the autonomy to accept, reject, or modify the outputs [32]. A broader vision has been presented in which copilots are viewed as collaborative partners that operate alongside human users, each enhancing the other’s capabilities [33]. This metaphor has further been extended to emphasize fluid, conversational interaction and limited autonomy in pursuing shared goals under human supervision [34]. Taken together, these perspectives highlight AI copilots as interactive, task-oriented systems that support users, particularly knowledge workers, through real-time assistance, domain adaptation, and human-guided decision-making. These shared characteristics form the foundation for a unified understanding of AI copilots presented in this survey.

### v2

@@ Abstract
AI copilots represent a new generation of AI-powered systems designed to assist users, particularly knowledge workers and developers, in complex, context-rich tasks. As these systems become more embedded in daily workflows, personalization has emerged as a critical factor for improving usability, effectiveness, and user satisfaction. Central to this personalization is preference optimization: the system’s ability to detect, interpret, and align with individual user preferences. While prior work in intelligent assistants and optimization algorithms is extensive, their intersection within AI copilots remains underexplored. This survey addresses that gap by examining how user preferences are operationalized in AI copilots. We investigate how preference signals are sourced, modeled across different interaction stages, and refined through feedback loops. Building on a comprehensive literature review, we define the concept of an AI copilot and introduce a taxonomy of preference optimization techniques across pre-, mid-, and post-interaction phases. Each technique is evaluated in terms of advantages, limitations, and design implications. By consolidating fragmented efforts across AI personalization, human-AI interaction, and language model adaptation, this work offers both a unified conceptual foundation and a practical design perspective for building user-aligned, persona-aware AI copilots that support end-to-end adaptability and deployment.
@@ 1 Introduction
AI-powered, context-aware systems—commonly referred to as AI copilots—are rapidly emerging as collaborative assistants for developers, analysts, and knowledge workers engaged in complex, high-context tasks. These systems are integrated into user workflows and designed to understand evolving goals, adapt to individual preferences, and provide real-time, personalized support. However, achieving meaningful personalization in such systems presents unique challenges. User intent is often ambiguous or changes over time, tasks may span multiple domains, and responses must be dynamically adapted to nuanced contexts. These systems must be capable of aligning with dynamic user preferences, as personalization is a key success factor for digital assistants, enabling them to adapt their responses and behavior to individual users’ routines and context [1]. At the core of this personalization lies preference optimization—the ability of a system to detect, interpret, and align with user preferences [2]. While advances in preference modeling have gained traction in machine learning and recommender systems, their integration into real-time, interactive systems like AI copilots remains limited and fragmented. Designing AI copilots that can continuously adapt to evolving preferences in high-context environments is still a major open challenge.
@@
Research related to this survey broadly falls into two categories: (i) AI systems and intelligent assistants, and (ii) preference optimization techniques. While each has been explored independently in the literature, their integration remains underexamined. The evolution of Virtual Assistants (VAs), such as Google Assistant, Siri, Cortana, and Alexa, has been well documented, with studies focusing on their core functionalities, user experience, and limitations in areas such as voice recognition, contextual awareness, and multi-turn dialogue management [3]. In enterprise contexts, Digital Assistants (DAs) have been examined through the lens of business integration, workflow support, and human-computer collaboration, highlighting both their organizational potential and associated risks [1]. Concerns around privacy and security in these systems have also been widely addressed, particularly in relation to data handling and unauthorized access [4]. Although the design and evaluation of AI copilots in retail has been recently analyzed through Microsoft’s case studies, the work is limited in scope and does not offer a comprehensive perspective on the broader cognitive assistant landscape [5]. On the other hand, preference optimization—especially in the context of aligning AI behavior with user intent—has seen significant theoretical and practical development. A recent comprehensive survey on Direct Preference Optimization (DPO) presents an in-depth analysis of its theoretical foundations, training methodologies, benchmark datasets, and diverse application scenarios [2]. However, such works primarily examine preference modeling in isolation and do not address its integration within interactive systems like AI copilots or digital assistants.
While existing studies contribute significantly to our understanding of intelligent assistant technologies and preference optimization methods, their integration in the context of AI copilots remains limited and underexplored. As a result, there is no consolidated view of how preference signals can be effectively incorporated into the behavior of real-time, context-aware systems. This survey aims to bridge these two streams of research by examining preference optimization techniques within the design of AI copilots—systems that operate collaboratively with users in high-context, evolving environments. Specifically, we address how preferences can be detected, modeled, and refined to enhance personalization and system responsiveness. The main contributions of this work are summarized as follows:
  • A unified and literature-grounded definition of an AI copilot is proposed, synthesizing diverse descriptions and terminology from recent studies across domains.
@@
We organized the content of our survey as follows: Section 2 provides foundational background by examining the historical development and evolving terminology associated with AI copilots. Section 3 introduces a unified, literature-informed definition of AI copilots, integrating insights from diverse fields. Section 4 presents a conceptual AI copilot architecture focused on preference optimization, offering a structured analysis of core components, including preference sources, detection techniques, personalization strategies, and feedback-driven refinement. Finally, Section 5 concludes with a summary of key insights and directions for future research.
@@ 2.1 AI Copilot Related Terms
2.1.1 Cognitive Assistant.
Cognitive computing has been characterized as the development of hardware and software systems that emulate human cognitive processes [6]. Within this paradigm, the concept of Metacognition has been introduced to extend these systems by enabling self-monitoring and adaptive reasoning, resulting in symbolic, closed-loop architectures where humans and machines exchange information in real-time feedback cycles [7]. A notable instantiation of this framework is the Joint Cognitive System, which conceptualizes human and machine agents as goal-driven collaborators dynamically sharing cognitive responsibilities within context-specific environments [8]. This approach shifts the focus from automation to cooperative cognition, emphasizing mutual adaptation between humans and intelligent systems. Cognitive assistants (CAs) have been positioned as key enablers of such collaborative paradigms. By merging human creativity with artificial intelligence’s computational capabilities, these systems have been shown to enhance problem-solving, decision-making, and situational awareness [9]. Studies in hybrid intelligence and empathic AI-supported interactions further highlight their potential to facilitate innovation, foster empathy, and support effective teamwork [10, 11]. This evolving synergy underscores the role of cognitive assistants in realizing adaptive and human-centered AI partnerships.
@@ 2.1.2 Digital Assistant
The term Digital Assistant (DA) is frequently interchanged with Virtual Assistant (VA), yet it represents a broader concept. In a study about its anatomy, Wells et al. [12] referred to it as a versatile tool and amorphous compound technology. Mobility, rapid data analysis capability, and supporting multiple interfaces are among its key features, making it ideal for task delegation and advice provision applications. A key feature of DAs has been identified as their capacity to combine intuitive natural language dialogue with personalized, context-specific assistance [13]. These systems have been categorized based on their modality and interaction style, with Smart Personal Assistants (SPAs) considered a personalized subset of DAs. A common classification framework distinguishes between text-based interfaces, often referred to as Chatbots, and voice-based systems, a categorization that has also been applied to DAs [14]. By considering Chatbots as a form of DA, all the categories of employed technologies in their construction can be categorized under DA. This allows DA’s boundaries to be expanded to cover a wider range of technologies, from simple rule-based Chatbots with limited capabilities to intelligent voice-based VAs with speech recognition capability [15].
Table 1: Comparison of AI copilot related terms.
@@
Cognitive Assistant Advanced decision support through reasoning and learning Proactive / Semi-autonomous Multimodal (text, voice, context-aware UI) Often domain-specific
(e.g., healthcare, law)
@@
Chatbot Text-based conversational responses Reactive Text Often domain-specific
(e.g., customer support, retail)
@@
AI Copilot Context-aware collaboration on complex or creative tasks Proactive Multimodal (text, code, UI, APIs) Task- or profession-specific
(e.g., software development, writing)
@@ 2.1.3 Virtual Assistant
As a good transitional definition from DA to VA, Apple Siri and Amazon Alexa as two popular intelligent assistants referred to as voice-based DAs [1, 16]. Their ability to perform hands-free operations, context understanding, and handle complex conversations are among the key features mentioned by Wellsandt et.al study [12]. Besides their speech recognition ability to understand voice commands and communicate over the same channel [4], voice-activation ability made them suitable for integrating into daily use devices [17]. Although they are mainly referred to as VAs, Smart Personal Assistant (SPA) is another term used for Apple Siri and Amazon Alexa as two device-integrated assistants focusing on their personalized assisting [13]. This positions them within the broader spectrum of DAs, primarily known for their voice-based interfaces, but also encompassing text-based interaction as an additional feature, distinguishing another subset of such assistants [18].
@@ 2.1.4 Chatbot
Chatbots are specialized programs tailored for customer service, helping users find desired information [19]. This focused definition underscores their application and emphasizes their role, setting them apart from other DAs despite shared technologies and capabilities. A defining feature of chatbots, as highlighted in literature [20], is simulating near-human conversations using natural language or text, to create an impression of engaging in dialogue with a human [21]. Emerging from the field of Human-Computer Interaction (HCI), which centres on enabling users to communicate with computers, Dialogue Systems is a term often used to describe Chatbots [19, 22]. Conversational Systems, Conversational Agents, and Chatter-bot are among other terms used interchangeably to refer to the same term [20].
While there are different criteria for chatbot classification in the literature, they are primarily classified by their applications, falling into either Task-oriented or Non-task-oriented (open-ended or simple chit-chat) categories [19, 21]. Task-oriented chatbots generally consist of domain-specific hand-crafted rules with limited and focused conversation context analysis and are typically utilized in devices for convenience. Non-task-oriented or informative chatbots often handle random dialogue contexts without aiming for a specific goal and may require online search capabilities to operate effectively [22, 21]. Based on the knowledge domain, chatbots can be categorized into closed-domain or open-domain, where the open-domain modeling, a challenging field, is predominantly addressed using Neural Network (NN) approaches, including both Retrieval-based and Generative (Generation-based) chatbots [20]. Retrieval-based chatbots support immediate response generation in e-business by ranking candidate responses based on similarity measurements between the query and the knowledge base of QA pairs. Data-driven generative chatbots utilize Natural Language Processing (NLP) to extract content from queries and learn from entire dialogue contexts, rather than solely focusing on the current turn. These chatbots apply both supervised methods, which rely on annotated data and manual feature extraction, and unsupervised methods, which automatically learn features from unlabeled data. Responses are generated word by word using neural network techniques such as RNN, Long Short-Term Memory (LSTM), and Deep Seq2Seq, all of which operate based on a learned vocabulary [20]. In this regard, Retrieval-Augmented Generation (RAG)-based approaches, as a more recent approach, combine the previous techniques for creating a better open-domain chatbot and even injecting more domain-specific knowledge by additional training [23].
@@ 2.1.5 Expert System
As one of the old terms in cognitive computing, an Expert System (ES) is defined as a computer program that emulates the behavior of a human expert [24]. They are a branch of AI (that is a high-level term) that is used for transferring task-specific knowledge from human to computer in a way that it continuously changes and obtains new knowledge[25]. Used in problem-oriented domains including social science, cognitive science, and human behaviour analysis, ES has two main parts that enable it to produce expert-like advice. As the first main component, the knowledge base, constructed by knowledge engineers using a suitable notation of language, stores the codified collected knowledge. The inference engine, guided by a suitable reasoning method chosen to ensure efficient use of the knowledge, then uses the knowledge to develop a conclusion [24]. For transparency reasons, the system can benefit from an explanation facility to elucidate the logic behind the conclusion [26]. Although Rule-Based Systems (RBS) or Knowledge-Based Systems (KBS) are used to refer to the same term [27, 26], others consider them as main streams of ES[25]. In the RBS, knowledge representation is a collection of IF-THEN rules. KBS, as a human-centred design, includes knowledge engineering tools and a user interface for this reason. Case-Based Reasoning (CBR) systems in one category of ESs adapt solutions to new challenges by comparing them to past cases based on set parameters. Meanwhile, Fuzzy ESs in another category, embrace uncertainty in human reasoning, transcending binary logic for nuanced decision-making in complex situations where traditional approaches fall short [25].
The Belief Rule-Based Expert System (BRBES) merges the accuracy of ESs with the adaptability of data-driven models, adeptly handling quantitative and qualitative data to address complexities in uncertain environments [28]. It manages uncertainties—like fuzziness, randomness, and ignorance—through belief degrees and employs the Evidence Reasoning (ER) algorithm for seamless data integration. This approach makes BRBES a transparent and dependable solution for decision-making in situations marked by scarce or ambiguous information.
@@ 3 AI Copilot Definition
AI copilots have been described as intelligent assistants designed to support users in executing cognitive tasks [5]. Their architectures are typically composed of large language models, mechanisms for knowledge retrieval and action execution, orchestration modules to manage subtasks, structured system prompts, and responsible AI safeguards to ensure alignment with human intent. In domain-specific applications such as pathology, copilots have been tailored through fine-tuning on real-world instructions to address complex tasks like question answering [29]. These systems are often multimodal and interactive, capable of processing both visual and textual inputs, and are used to assist human-in-the-loop decision-making processes. In programming contexts, copilots have been implemented to provide real-time suggestions that support knowledge workers and improve productivity, while allowing users the autonomy to accept, reject, or modify the outputs [30]. A broader vision has been presented in which copilots are viewed as collaborative partners that operate alongside human users, each enhancing the other’s capabilities [31]. This metaphor has further been extended to emphasize fluid, conversational interaction and limited autonomy in pursuing shared goals under human supervision [32].
Taken together, these perspectives highlight AI copilots as interactive, task-oriented systems that support knowledge workers through real-time assistance, domain adaptation, and human-guided decision-making. These shared characteristics form the foundation for a unified understanding of AI copilots presented in this survey.